# Extreme Conformal Prediction: Reliable Intervals for High-Impact Events

### v1

@@ 5.1 Description and aim
Flooding is one of the most impactful natural hazards in terms of infrastructure and economic damage, and of the endangerment of human lives. Methods from extreme value theory have proven successful for assessing flood risk and providing reliable worst-case scenarios (e.g., Katz et al., 2002; Keef et al., 2009; Asadi et al., 2018; Engelke and Hitz, 2020; Engelke and Ivanovs, 2021).
@@
In Switzerland, the Federal Office for the Environment (FOEN) monitors the river flow with numerous gauging stations throughout the river network. In its capital city, Bern, extreme water flow events of the Aare river led to several major floods, causing some of the most severe infrastructural and economic flooding damages recorded in the country. The main driver of strong water-flow events is the cumulative amount of upstream precipitation. In this study, we rely on the average daily discharge measures of the Aare river (in $\text{m}^{3}\text{s}^{-1}$) provided by the FOEN [^1] https://www.hydrodaten.admin.ch/., and on recordings of daily precipitation (in mm) at various meteorological stations, obtained from MeteoSwiss [^2] https://gate.meteoswiss.ch/idaweb.. This version of the dataset was preprocessed and analysed in previous studies (Pasche et al., 2023; Pasche and Engelke, 2024).
@@
With our proposed extreme conformal approach, we aim to provide high-confidence one-day-ahead interval forecasts of the conditional range for water flow. We rely on several extreme quantile regression models pretrained to forecast the one-day-ahead extreme quantiles of the Aare water flow in Bern, given observations of upstream precipitation at six locations in Bern’s water catchment and of the average daily water flow at an upstream gauging station, during the previous 10 days. Figure A.2 in Appendix A.1 shows the location of those meteorological and gauging stations. We then use the GPD safeprofile extreme conformalization approach recommended in Section 4 to obtain the desired one-day-ahead PIs. The extreme quantile regression models we consider are EQRN, GBEX (Velthoen et al., 2023), EGAM (Youngman, 2019), and EXQAR (Li and Wang, 2019). We also consider the simple unconditional GPD quantile estimates as a comparative baseline.
@@
The models were pretrained on data from 1939 to 1951. They were all fine-tuned with a grid search for hyperparameter selection. We use data from 1958 to 1999 for calibration and testing (48 years). The observations after 1999 are not considered, due to a major distribution shift [^3] See the flood report of the FOEN at https://www.hydrodaten.admin.ch/en/2135.html.. We choose 10 years as the default calibration size in the first part of the analysis, but vary this size from 3 to 15 years in the second part. We use multiples of complete years to keep a seasonal balance in the calibration and test sets.
@@ 5.2 Results
Figure 3 compares the number of observations exceeding the PIs from each method during the test period, using predictions of the different pretrained models, for a range of moderate to extreme confidence levels. The number of observations expected to exceed the PIs during the 38-year test period varies from 2,776, for $0.81-\alpha=$ 0.8, to only $1.4$, for the largest level $0.99991-\alpha=$ 0.9999. Using the original model predictions as PIs leads, in most cases, to undercoverage. Although the best-performing predictions seem to vary around the target coverage, they fail to provide satisfactory coverage consistently. The classical conformalization seems effective for the lowest two confidence levels but worsens the coverage for the following moderately extreme levels, compared to the initial predictions. At the largest level, going beyond the calibration order statistics, the classical method yields infinite-endpoint PIs. The extreme conformalization method yields finite PIs with significantly better coverage for confidence levels above $0.95$, for which the approaches differ. The PI coverage consistently strictly satisfies the target confidence levels for each initial prediction model. The profile likelihood procedure is stable for almost all models and confidence levels. It was replaced by the bootstrap PIs only once, with the EGAM predictions.
@@
Figure 4 shows the initial EQRN predictions, which seem to fit the data variation best, and their conformalized PI endpoints for two of the considered confidence levels, including the largest. The infinite classical PIs are non-informative for the largest level. On the other hand, the extreme PIs, providing correct test coverage, are not overly large compared to the data variation, the original predictions, and the unconditional quantile estimates. All other extreme conformal corrections for the EQRN predictions are even smaller in magnitude. During the considered year, the original predictions at the largest confidence level are exceeded once, on 25th July 1973. This exceedance is covered by the conformalized extreme PI.
Figure 5 shows the evolution of the test coverage with the calibration size, for all predictions, methods, and confidence levels. We observe that the extreme PIs significantly outperform the classical conformalization in terms of empirical test coverage, for all relevant levels. It always provides informative finite PIs, contrary to the classical approach that yields infinite PIs when $\alpha<1/(n_{c}+1)$, that is, for calibration sizes up to 8 years with $0.99971-\alpha=$ 0.9997, and for all sizes with $0.99991-\alpha=$ 0.9999. The extreme PI has valid coverage in almost all combinations, including when the classical approach and/or original predictions significantly undercover. In the few undercovered situations, it still outperforms both the original predictions and the classical approach.
Figure 5: Empirical test coverage of the prediction intervals for a range of calibration-set sizes, for each conformalization method, pretrained model (columns), and confidence level $1-\alpha$ (rows). For $0.951-\alpha<$ 0.95, corresponding to the quantile level of the GPD threshold $u$, the GPD approach coincides with the classical.
@@ 6 Conclusion
We propose a conformalization method which relies on extreme value statistics to provide conservative and non-degenerate prediction intervals for reliable risk assessment of high-impact events. The novel method uses the well-studied peaks-over-threshold approach, which leverages the generalized Pareto distribution to extrapolate the necessary conformal correction on the calibration data to the required extreme confidence levels. It uses a conservative confidence interval solution for robustness against estimation and approximation biases.
@@
In the simulation study and the application to river-flow forecasts, our extreme conformal prediction method consistently provides better coverage than both the original predictions and the classical conformalization method at large confidence levels. For the largest levels, the classical approach results in infinite (or undefined) PIs, which are of no practical use. On the other hand, our recommended method yields informative finite intervals that consistently achieve the desired coverage for confidence levels up to several orders of magnitude larger than what is feasible with classical methods. Importantly, our extreme conformal prediction approach can be used in combination with any extreme quantile regression model, including black-box machine learning methods without known asymptotic guarantees.
@@
One downside of our approach is its potential overconservativeness in certain scenarios, e.g. for moderately extreme confidence levels and some lighter-tailed data distributions. This is, at least in part, a consequence of our conservative CI-based solution, used to circumvent possible estimation and/or approximation biases in the peaks-over-threshold GPD estimation. Contrary to the finite sample properties of classical conformal prediction, the GPD relies on asymptotic results and lacks those finite sample guarantees. This is an unavoidable trade-off for the ability to extrapolate beyond the moderate levels for which the classical empirical-quantile-based method is feasible. In the general framing of this work, strict finite-sample guarantees for quantiles extrapolated beyond the data range are not theoretically possible without additional assumptions, due to the broad spectrum of potential sub-asymptotic tail behaviours (Thomas, 2015; Lhaut et al., 2022).
@@
This work establishes a first method for extreme-confidence PIs in the context of the well-established split-conformalized regression. Further work is possible to include more specialised or alternative conformal approaches. One direction is to enhance the efficiency of data usage, using, for example, full conformal prediction or potential variants of jackknife (Barber et al., 2021; Steinberger and Leeb, 2016), infinitesimal jackknife (Alaa and van der Schaar, 2020), or cross-validation (Vovk, 2015). However, such approaches typically have a much greater computational cost, as they require refitting the predictive model multiple times. Other directions include extensions to non-stationary data (see e.g. Barber et al., 2023). Finally, approaches other than conformalization could also warrant investigations, such as building PIs based on the so-called high-quality criterion and using deep learning (Pearce et al., 2018; Khosravi et al., 2011). How to incorporate extreme value statistics to extrapolate prediction intervals to cover high-impact events in these methodologies appears to be largely open.
@@ Acknowledgements
In the Application to flood risk forecasting, we use river discharge and precipitation data recorded in Switzerland between 1930 and 1999, in the Rhine and Aare basins. The precipitation records can be ordered online for free from MeteoSwiss, on https://gate.meteoswiss.ch/idaweb, and the discharge records from the Swiss Federal Office for the Environment (FOEN), on https://www.bafu.admin.ch/bafu/en/home/topics/water/state/data/obtaining-monitoring-data-on-the-topic-of-water/hydrological-data-service-for-watercourses-and-lakes.html.
@@ Appendix A.1 Additional figures
A.1.1 Simulation study
Figure A.1 shows the distribution of the computed test coverage for each considered conformalization method, confidence level, and calibration size, for the light-tailed Gaussian-noise data and ground truth original predictions.

### v2

@@ 5.1 Description and aim
Flooding is one of the most impactful natural hazards in terms of infrastructure and economic damage, and of the endangerment of human lives. Methods from extreme value theory have proven successful for assessing flood risk and providing reliable worst-case scenarios (e.g., kat2002; kee2009; ASADI2018182; eng2018; Engelke2020).
@@
In Switzerland, the Federal Office for the Environment (FOEN) monitors the river flow with numerous gauging stations throughout the river network. In its capital city, Bern, extreme water flow events of the Aare river led to several major floods, causing some of the most severe infrastructural and economic flooding damages recorded in the country. The main driver of strong water-flow events is the cumulative amount of upstream precipitation. In this study, we rely on the average daily discharge measures of the Aare river (in $\text{m}^{3}\text{s}^{-1}$) provided by the FOEN [^1] https://www.hydrodaten.admin.ch/., and on recordings of daily precipitation (in mm) at various meteorological stations, obtained from MeteoSwiss [^2] https://opendatadocs.meteoswiss.ch/.. This version of the dataset was preprocessed and analysed in previous studies (Pasche2021; Pasche2024).
@@
With our proposed extreme conformal approach, we aim to provide high-confidence one-day-ahead interval forecasts of the conditional range for water flow. We rely on several extreme quantile regression models pretrained to forecast the one-day-ahead extreme quantiles of the Aare water flow in Bern, given observations of upstream precipitation at six locations in Bern’s water catchment and of the average daily water flow at an upstream gauging station, during the previous 10 days. Figure˜A.3 in Appendix˜A.2 shows the location of those meteorological and gauging stations.
5.2 Methodology
The river data exhibits a strong seasonal pattern with more water flow, and with more frequent and intense extreme events, during the late spring and summer months, due to snowmelt and heavy precipitation. This seasonality is likely to propagate to the residual nonconformity scores, violating the exchangeability assumption underlying classical conformal prediction. To account for it, we use the weighted variation of our extreme conformalization approach, discussed in Section˜3.3. As a general procedure to account for distribution drift, Barber2023nonstatConform suggest choosing large weights for recent calibration observations, close in time to the test period, and having weights decay for earlier observations, either exponentially or as a cutoff. Here, we make use of the inherent periodicity of the seasonal river discharge behaviour, by matching it with sinusoidal weights. Each year is partitioned into 24 roughly equal seasonal blocks of 15 to 16 days each. Let $B(i)\in\{1,\ldots,24\}$ denote the block in which an observation indexed by $i$ falls into. For each block $b=1,\ldots,24$, $\hat{q}^{e}_{\alpha}(b)$ is estimated as the upper-endpoint of the appropriate extreme score quantile CI, as described in Section˜3.2, but using the weighted GPD likelihood in Eq.˜15, with calibration sample weights
$w_{i}=\cos\left[\frac{2\pi}{24}\{B(i)-b\}\right]+1,\quad i=1,\ldots,n_{c}.$ (17)
This choice of weights gives the highest importance to calibration observations in the same seasonal block $b$ as the test point, and decreases the importance of observations in blocks further away in the year, with a yearly periodicity. For the estimation of $\hat{q}^{e}_{\alpha}(b)$, we use GPD safeprofile CI procedure recommended in Section˜4, although almost no numerical issues were encountered with the profile likelihood.
Our final one-day-ahead extreme PI for the average daily discharge, during seasonal block $b=1,\ldots,24$ of the year, given past observations of upstream precipitation and water flow $\bm{X}_{\text{test}}=\bm{x}$, is then $\hat{C}^{e}(\bm{x})=\left(0,\;\hat{Q}^{e}_{1-\alpha}(\bm{x})+\hat{q}^{e}_{\alpha}(b)\right]$. We compare this PI to the unconformalized quantile predictions $\hat{Q}^{e}_{1-\alpha}(\bm{x})$, and to a single-sided (see Section˜3.1) and weighted (Barber2023nonstatConform) variation of the classical split conformalized quantile regression PI $\hat{C}(\bm{x})$ (ConformalizedQR), using the same sinusoidal weights from Eq.˜17.
Apart from its seasonality, the data is also sequentially dependent. However, the residual dependence between the scores $S_{i}$ is likely weak and short-term, which should not significantly affect the marginal coverage guarantee of the conformal PIs (OliveiraConformNonexchError).
We consider several choices of extreme quantile regression models for $\hat{Q}^{e}_{1-\alpha}(\cdot)$: EQRN, GBEX (gbex), EGAM (ExGAM2), and EXQAR (EXQAR). We also consider the constant unconditional GPD quantile estimates as a comparative baseline. We emphasise results using the recurrent version of EQRN, which is specifically designed for sequential dependence, and seems to fit the data best (Pasche2024).
@@ 5.2 Methodology
The quantile models were pretrained on data from 1939 to 1951. They were all fine-tuned with a grid search for hyperparameter selection. We use data from 1951 to 1999 for calibration and testing, i.e. 48 years of daily data. The observations after 1999 are not considered, due to a major distribution shift [^3] See the flood report of the FOEN at https://www.hydrodaten.admin.ch/en/2135.html.. We choose the first 10 years as the default calibration set in the first part of the analysis, but vary its size from 3 to 15 years in the second part, and use the rest for estimating PI coverage empirically. We use multiples of complete years to keep a seasonal balance in the calibration and test sets.
@@ 5.3 Results
Figure˜3 compares the number of observations exceeding the PIs from each method during the test period, using predictions of the different pretrained models, for a range of moderate to extreme confidence levels. The number of observations expected to exceed the PIs during the 38-year test period varies from 2,776, for $1-\alpha=0.8$, to only $1.4$, for the largest level $1-\alpha=0.9999$. Using the original model predictions as PIs leads, in most cases, to undercoverage. Although the best-performing predictions seem to vary around the target coverage, they fail to provide satisfactory coverage consistently. The classical conformalization seems effective for the lowest two confidence levels but worsens the coverage for the following moderately extreme levels, compared to the initial predictions. At the two largest levels, the classical method yields uninformative infinite PIs. The extreme conformalization method yields finite PIs with significantly better coverage for confidence levels above $0.95$, for which the approaches differ. The PI coverage consistently strictly satisfies the target confidence levels for each initial prediction model. The profile likelihood procedure was stable for almost all models, confidence levels and seasonal blocks. A numerical instability only occurred once, for the unconditional predictions at the largest confidence level, during a single one of the seasonal blocks.
@@
Figure˜4 shows the initial EQRN predictions, which seem to fit the data variation best, and their conformalized PI endpoints for two of the considered confidence levels, including the largest, during a test year. At $1-\alpha=0.999$, even though it has more marginal coverage, the extreme PIs are not excessively larger than the classical PIs. The extreme intervals are even tighter during some of the seasonal blocks. At the largest level, the infinite classical-method PIs are uninformative. On the other hand, the extreme PIs, satisfying the desired test coverage, are again not overly large compared to the data variation, the original predictions, and the unconditional quantile estimates. Extreme conformal corrections for the EQRN predictions at the other unshown levels are smaller in magnitude. During the considered year, the original predictions at the largest confidence level are exceeded once, on 25th July 1973. This exceedance is covered by the conformalized extreme PI. Finally, we observed the localised adaptivity of weighted conformalization, as, for example, the conformal correction is larger during late summer than during winter for both methods and confidence levels.
Figure˜5 shows the evolution of the test coverage with the calibration size, for all predictions, methods, and different confidence levels. We observe that the extreme PIs significantly outperform the classical conformalization in terms of empirical test coverage, for all relevant levels. It always provides informative finite PIs, contrary to the classical approach that yields infinite PIs for calibration sizes up to 5 years with $1-\alpha=0.999$, and for all sizes at the two largest levels. The extreme PI has valid coverage in almost all combinations, including when the classical approach and/or original predictions significantly undercover. In the few undercovered situations, its coverage is closer to the target $1-\alpha$ than both the original predictions and the classical approach. For some of the lower levels, there seems to be a pattern of decreasing coverage with increasing calibration size for the conformalized PIs. The largest discharge events happening in years one, three and 15 of the chosen calibration period is a likely explanation for this decrease in between, although there might also be a small effect from the decreasing test size, as observed with the coverage of the unconformalized predictions.
Figure 5: Empirical test coverage of the prediction intervals for a range of calibration-set sizes, for each conformalization method, pretrained model (columns), and confidence level $1-\alpha$ (rows). For $1-\alpha<0.95$, corresponding to the quantile level of the GPD threshold $u$, the GPD approach coincides with the classical.
@@ 6 Conclusion
We propose a conformalization method which relies on extreme value statistics to provide conservative and nondegenerate prediction intervals for reliable risk assessment of high-impact events under extreme-confidence requirements. The novel method uses the well-studied peaks-over-threshold approach, which leverages the generalized Pareto distribution to extrapolate the necessary conformal correction from the calibration data to the required extreme confidence levels. It uses a conservative confidence interval solution for robustness against estimation and approximation biases.
@@
In the simulation study and the application to forecasting river flow, at large confidence levels, our extreme conformal prediction method consistently provides PIs with significantly better coverage properties than both the original predictions and classical conformalization. For the largest levels, classical approaches result in infinite (or undefined) PIs, which are of no practical use. On the other hand, our recommended method yields informative finite intervals that consistently achieve the desired coverage for confidence levels up to several orders of magnitude larger than what is feasible with classical methods. The weighted version of our approach can account for nonstationary data drifts, such as the seasonal behaviour of the river flow. Importantly, our extreme conformal prediction method can be used in combination with any extreme quantile regression model, including black-box machine learning methods without known asymptotic guarantees.
@@
One downside of our approach is its potential overconservativeness in certain scenarios, e.g. for moderately extreme confidence levels and some lighter-tailed data distributions. This is, at least in part, a consequence of our conservative CI-based solution, used to circumvent possible estimation and approximation biases in the GPD estimation. Moreover, peaks-over-threshold approaches rely on asymptotic properties. Thus, our approach lacks the finite-sample guarantees of classical conformal prediction. However, this is an unavoidable trade-off for the ability to extrapolate beyond the moderate levels for which the classical empirical-quantile-based method is feasible. In fact, obtaining finite-sample bounds for quantiles extrapolated beyond the data range is not theoretically possible without additional assumptions on the data distribution (BoucheronThomasConcentration; ThomasThesisEVTConcentr; LhautEVTConcentr). Another example of such an unavoidable trade-off in conformal prediction is the impossibility of achieving conditional coverage in finite samples, which led to asymptotic conditional coverage as an alternative guarantee (LeiWassermanConformalLocalAsymptotic).
@@
This work establishes a first method for extreme-confidence PIs, as an extension of the well-established conformal regression framework. As our experiments mainly focused on the split-conformal approach, due to its computational efficiency and practical popularity, further work is possible with alternative or more specialised conformal approaches. One direction is to enhance the efficiency of data usage, using, for example, the full conformal version of our approach, described in Appendix˜A.1, or potential variants of jackknife (barber2021predictive; steinberger2016leave), infinitesimal jackknife (alaa2020discriminative), or cross-validation (vovk2015cross). However, such approaches typically have a much greater computational cost, as they require refitting the predictive model many times. Whether the gain in statistical efficiency for our extreme conformal approach from using these alternative conformalization procedures is worth the extra computational cost remains an open question. Other directions include further extensions to nonexchangeable data. The weighted version of our extreme conformal approach seems suitable for nonstationary covariate drifts and shifts (analogously to Barber2023nonstatConform; TibshiraniCoformShift). However, alternative solutions might be necessary to account for other problematic scenarios, such as long-term dependence or drifts in the conditional distribution of the response given the covariates. Finally, approaches other than conformalization could also warrant investigations, such as building PIs based on the so-called high-quality criterion and using deep learning (DLpredintervals; khosravi2011comprehensive; ChenLam21PI). How to incorporate extreme value statistics to extrapolate prediction intervals to cover high-impact events in these methodologies appears to be largely open.
@@ Acknowledgements
In the Application to flood risk forecasting, we use river discharge and precipitation data recorded in Switzerland between 1930 and 1999, in the Rhine and Aare basins. The precipitation records can be freely obtained from MeteoSwiss, on https://opendatadocs.meteoswiss.ch/, and the discharge records from the Swiss Federal Office for the Environment (FOEN), on https://www.bafu.admin.ch/bafu/en/home/topics/water/data-and-maps/water-monitoring-data/hydrological-data-service-for-watercourses-and-lakes.html.
Implementation and reproducibility
An open-source R package implementing the proposed extreme conformal procedure, including its weighted variant for nonstationary data, will be released at https://github.com/opasche/ExtremeConformal. Its development required an efficient and versatile algorithm for estimating profile-likelihood confidence intervals for extreme quantiles (and return levels), which will be released as a separate R package at https://github.com/opasche/ExtremeCI. The code and data, with detailed instructions for reproducing the results presented in this paper, will also be made available at https://github.com/opasche/Reprod_ExtremeConformalPred.
@@
Appendix A.1 Extensions to other conformal approaches
This section describes in more detail the extensions of our proposed extreme conformal method to alternative conformal procedures discussed in Section˜3.3.
As explained in the main text, the extension to different base predictive models and scores is the most straightforward, as our extreme conformalization described in Section˜3.2 is agnostic to their definition. This includes the classical split-conformal approach that, instead of the quantile predictions $\hat{Q}_{\alpha/2}(\bm{x}),\hat{Q}_{1-\alpha/2}(\bm{x})$, relies on a conditional-mean base regression model $\hat{\mu}(\bm{x})$, and the residuals $s(\bm{x},y):=\left\lvert y-\hat{\mu}(\bm{x})\right\rvert$ as nonconformity score (SplitConformalMargProof; papadopoulos2008inductive; Lei18ConformalPred). The procedure from Section˜3.2 would then be performed with these alternative scores, resulting in the fixed-length extreme conformal PIs $\hat{C}^{e}(\bm{x})=\left[\hat{\mu}(\bm{x})-\hat{q}^{e}_{\alpha},\hat{\mu}(\bm{x})+\hat{q}^{e}_{\alpha}\right]$. If a residual-dispersion estimate $\hat{\sigma}(\bm{x})$ is also available, for instance, from a heteroscedastic regression model or a Bayesian approach, using the scaled residuals $s(\bm{x},y):=\left\lvert y-\hat{\mu}(\bm{x})\right\rvert/\hat{\sigma}(\bm{x})$ as nonconformity scores (papadopoulos2008normalizedCoform; papadopoulos2011regressionknn; Lei18ConformalPred) would result in the varying-length extreme PIs $\hat{C}^{e}(\bm{x})=\left[\hat{\mu}(\bm{x})-\hat{q}^{e}_{\alpha}\hat{\sigma}(\bm{x}),\hat{\mu}(\bm{x})+\hat{q}^{e}_{\alpha}\hat{\sigma}(\bm{x})\right]$. Although the latter doesn’t require extreme quantile regression to yield varying-length PIs, it tends to underestimate conditional variability and yield less adaptive predictions (ConformalizedQR).
In full-conformal procedures, the data is not split into training and calibration sets. Instead, given a training set $\{(\bm{X}_{i},Y_{i})\}_{i=1}^{n}$ and the test covariates $\bm{X}_{\text{test}}$, the PI $\hat{C}(\bm{X}_{\text{test}})$ for $Y_{\text{test}}$ is constructed by refitting the base model $\hat{f}$ (e.g., $\hat{Q}_{1-\alpha}$ or $\hat{\mu}$) on $\{(\bm{X}_{i},Y_{i})\}_{i=1}^{n}\cup\{(\bm{X}_{\text{test}},y)\}$, for a dense grid of values $y$ in the value space of $Y$. Each fit is denoted by $\hat{f}^{y}$. The desired nonconformity scores are then defined, for each $y$, as $S_{i}(y):=s(\bm{X}_{i},Y_{i};\hat{f}^{y})$, $i=1,\ldots,n$ (vovk1999ml; vovk2005algorithmic; shafervovk2008tutorial). Consequently, for a full-conformal variant, our proposed procedure from Section˜3.2 should be performed, for each $y$, on the scores $\{S_{i}(y)\}_{i=1}^{n}$, to obtain a $\hat{q}^{e}_{\alpha}(y)$. The resulting PI is then $\hat{C}^{e}(\bm{X}_{\text{test}})=\{y:s(\bm{X}_{\text{test}},y;\hat{f}^{y})\leq\hat{q}^{e}_{\alpha}(y)\}$. Although the latter makes a more efficient use of the data than the split variant, it is extremely computationally costly. The analogous extension to the middle-ground $k$-fold approaches, such as Jackknife+/CV+ (barber2021predictive) and cross-conformal prediction (vovk2015cross), would also suffer, less extremely, from similar refitting expensiveness.
@@ Appendix A.2 Additional figures
A.2.1 Simulation study
Figure˜A.1 shows the distribution of the computed test coverage for each considered conformalization method, confidence level, and calibration size, for the light-tailed Gaussian-noise data and the ground truth as base predictions. Figure˜A.2 shows the same test-coverage distribution, for each considered conformalization method, confidence level, and calibration size, for the Student $t$ distributed noise and the linear GPD quantiles as the base predictions.